Find coefficient of correlation of the following data. Use this data to estimate y when x = 12


X 3 6 7 10


Y 12 18 25 22






1
Expert's answer
2021-07-06T12:53:20-0400

By using Excel function =CORREL(A2:A5,B2:B5)



The coefficient of correlation = 0.7910

y=a + bx

=INTERCEPT(B2:B5,A2:A5)

a = 9.24

=SLOPE(B2:B5,A2:A5)

b= 1.54

y=9.24+1.54x

When x=12

"y(12) = 9.24 + 1.54 \times 12 = 27.72"